# Disambiguation for Video Frame Interpolation

### v3

@@ VI-D User study
VI-F Multi-frame qualitative comparison
As shown in Fig. 10, we present a qualitative comparison of different variants of each VFI model under multi-frame setup. $[D,M]_{e}$ denotes the model that incorporate both the indexing map with multi-frame refiner and continuous indexing map estimator. As anticipated, the $[D,M]_{e}$ configuration consistently yields the highest visual quality across all base models, demonstrating the effectiveness of both the multi-frame refiner and the continuous indexing map estimator.
Figure 10: Qualitative comparison under multi-frame setting. $[T]$: original arbitrary time VFI models using time indexing. $[D]_{u}$: models trained using distance indexing, then inference using uniform maps. $[D,R]_{u}$: models trained using both distance indexing and iterative reference-based estimation, then inference using uniform maps. $[D,M]_{e}$: models trained using distance indexing with multi-frame fusion, then inference using estimated maps.
VI-G Multi-frame quantitative comparison
In Table VII, we present a quantitative performance comparison across various model variants. The configurations $[T]$, $[D]_{u}$, and $[D]$ retain the same definitions as previously described. $[D]_{e}$ denotes the model augmented with a continuous indexing map estimator, $[D,M]$ includes the proposed multi-frame fusion with the refiner model, and $[D,M]_{e}$ incorporates both the refiner and the map estimator. On pixel-level metrics such as PSNR and SSIM, the enhanced variant $[D,M]_{e}$ consistently outperforms the base model $[T]$ as well as $[D]_{u}$, which lacks a map estimator and relies on uniform indexing maps. The improvement from $[D]_{u}$ to $[D,M]_{e}$ highlights the effectiveness of integrating the continuous indexing map estimator. Meanwhile, performance gains from $[D]$ to $[D,M]$, and from $[D]_{e}$ to $[D,M]_{e}$, demonstrate the additional benefit provided by the multi-frame refiner.
TABLE VII: Multi-frame comparison on Vimeo90K septuplet dataset. $[M]$ denotes the multi-frame fusion. $[\cdot]_{e}$ denotes inference with estimated indexing map. Other notations have the same meaning as in the previous experiments.
RIFE [7] $[T]$ $[D]_{u}$ $[D]_{e}$ $[D,M]_{e}$ $[D]$ $[D,M]$
PSNR $\uparrow$ 28.22 27.55 28.25
28.34
29.20
@@ VI-G Multi-frame quantitative comparison
IFRNet [24] $[T]$ $[D]_{u}$ $[D]_{e}$ $[D,M]_{e}$ $[D]$ $[D,M]$
PSNR $\uparrow$ 28.26 27.40 27.63
28.28
29.25
@@
VI-H Multi-frame tuning strategy
In this section, we evaluate different training strategies for multi-frame fusion using RIFE [7] as the baseline for ablation studies. All experiments follow the same setting of previous experiments on Vimeo90k. We find that jointly tuning both refiner and map estimator leads to instability and failure to converge. Thus, we split the tuning process in two stages. In the first stage, we first finetune the multi-frame refiner $\mathcal{F^{\prime}}$ alone using original VFI loss. In the second stage, there are two possible choices. Since jointly tuning with the map estimator remains unstable, we either finetune the flow estimator only to adapt it to VFI (referred to as FE), or finetune the VFI modules $\mathcal{F}$ and $\mathcal{F^{\prime}}$ to adapt them to the pretrained flow estimator (referred to as FE), as shown in Table VIII. The loss term $\mathcal{L}_{D}$ is calculated by $\mathcal{L}_{D}=\|D_{t}-D_{t}^{\text{RAFT}}\|_{2}^{2}$, denoting the L2 loss between the indexing maps from CPFlow and RAFT. The setting $\mathcal{L}_{V_{f}}$ corresponds to training only the VFI refiner with its original loss, while $\mathcal{L}_{V_{b+f}}$ involves jointly training both the base VFI model and the refiner with their original loss. Among the tested strategies, jointly tuning both the VFI model and the refiner under $\mathcal{L}_{V_{b+f}}$ achieves the best performance. This improvement is attributed to the model’s ability to adapt both components to the updated indexing map derived from the newly predicted optical flow.
TABLE VIII: Multi-frame ablation study
@@ VI-H Multi-frame tuning strategy
TABLE IX: Costs comparison including runtime and number of parameters. We evaluate the computational overhead on an NVIDIA A100 GPU using images with a resolution of 448×256.
$[D]$ $[D,M]$ $[D,M]_{e}$
@@
VI-I Computational costs of the proposed framework
@@
The runtime and the number of parameters are reported in Table IX. The estimation process introduces an additional latency of approximately 0.04 seconds per inference, which is relatively low. The number of parameters for the map estimator is 10.2M. Overall, the computational overheads introduced by indexing map estimation are acceptable in practical scenarios.
@@
We also evaluate the computational overhead introduced by the multi-frame refiner as reported in Table IX. The refiner adds approximately 0.03 seconds per inference and requires an additional 10.3 million parameters, which is practical for deployment.
VII Conclusion and Future Work
@@ Acknowledgments
This work was partially supported by the National Key R&D Program of China (No. 2022ZD0160104). We thank Dorian Chan, Zhirong Wu, and Stephen Lin for their insightful feedback and advice. Our thanks also go to Vu An Tran for developing the web application.
Appendix A Proof of Eq. 4
Eq. 4 in the paper can be rigorously proven if an L2 loss is used,
@@ Appendix B The Rationale for Solving Ambiguity
We claim the “correct” distance map is hard to estimate accurately from merely two frames since there are a wide range of possible velocities. If considering more neighboring frames like Xu et al. [8] (more observation information), it is possible to estimate an accurate distance map for pixel-wise aligned interpolation, which we leave for future work.
@@
TABLE X: Comparison on Vimeo90K triplet dataset. $[T]$ denotes the method trained with traditional fixed time indexing paradigm. $[D]$ denotes the distance indexing paradigm. $[\cdot]_{u}$ denotes inference with uniform map as time indexes.
@@ Appendix C Additional Experiments and Analysis
While the benefits of our proposed disambiguation strategies are best demonstrated on arbitrary-time VFI models, they actually improve the performance of fixed-time models as well. Using RIFE [7] as a representative example, we extend our comparison to the fixed-time training paradigm, depicted in Fig. 11. The label $[T]$ RIFE (Tri) refers to the model trained on the triplet dataset from Vimeo90K [54] employing time indexing. Conversely, $[D]$ RIFE (Tri) indicates training on the same triplet dataset but utilizing our distance indexing approach. Both $[D]$ RIFE and $[D,R]$ RIFE models are trained on the septuplet dataset, consistent with our earlier comparison. Despite being trained on varied datasets, it is evident that the arbitrary time model outperforms the fixed time model. However, the efficacy of distance indexing appears restrained within the fixed-time training paradigm. This limitation stems from the fact that deriving distance representation solely from the middle frame yields a sparse distribution, making it challenging for the network to grasp the nuances of distance. We delve deeper into the quantitative analysis of these findings in Table X. Compared to training arbitrary time models on the septuplet dataset, the advantages of distance indexing become notably decreased when training fixed time models on the triplet dataset.
TABLE XI: Comparison on the septuplet of Vimeo90K [54] using LPIPS loss [56]. We use RIFE [7] as a representative example.
$[T]$ $[D]_{u}$ $[D,R]_{u}$
PSNR $\uparrow$
27.19
26.71 26.72
SSIM $\uparrow$
0.898
0.889 0.890
LPIPS $\downarrow$
0.061
0.065 0.064
NIQE $\downarrow$ 6.307 5.901
5.837
Comparison of using perceptual loss
In addition to training with traditional pixel losses based on L1 and L2 losses, we present the results of employing the more recent LPIPS loss [56] with a VGG backbone [61], as shown in Table XI. The non-reference perceptual quality metric, NIQE, shows notable improvement across all variants. The results also consistently demonstrate the effectiveness of our strategies in resolving velocity ambiguity. Besides, due to the direct optimization of LPIPS loss, the assumption of constant speed in uniform maps affects the performance for this metric. This is why in the test results, $[T]$ has a lower LPIPS, while $[D]_{u}$ and $[D,R]_{u}$ are slightly higher.
TABLE XII: Comparison with $[D_{x,y}]$ scheme on Vimeo90K septuplet dataset. $[D_{x,y}]$ denotes training with two channel distance map with $\bm{x}$ and $\bm{y}$ directions. $[\cdot]_{u}$ denotes inference with uniform maps.
$[T]$ $[D]_{u}$ $[D,R]_{u}$ $[D_{x,y}]_{u}$ $[D_{x,y},R]_{u}$
@@
Figure 12: User interface of user study.
Two channel scheme
Comparison results with the two channel scheme, i.e., $D_{t}$ with $\bm{x}$ and $\bm{y}$ directions, are shown in Table XII. The observations are as follows: (1) $[D_{x,y}]_{u}$ outperforms $[D]_{u}$, which makes sense since it accounts for both speed and direction. (2) $[D,R]_{u}$ performs better than $[D_{x,y}]_{u}$. Our speculation is that $[R]$ benefits not only from addressing the directional ambiguity but also from reducing the prediction difficulty via divide-and-conquer. (3) $[D_{x,y},R]_{u}$ does not exceed $[D,R]_{u}$, showing that the iterative formulation is sufficient to resolve the directional ambiguity. $[D_{x,y}]_{u}$ has the potential to learn different trajectories but cannot realize that potential since the trajectory distribution within a short time is not sufficiently diverse. However, $[D_{x,y}]_{u}$ only involves increasing the input channels and does not need to run iteratively. Thus, $[D_{x,y}]_{u}$ is a better choice for fast/lightweight frame interpolation.
@@
Appendix D User Study UI

### v4

@@ VI-E Evaluating distance indexing on multi-frame baseline
VI-G Multi-frame qualitative comparison
As shown in Fig. 12, we present a qualitative comparison of different variants of each VFI model under multi-frame setup. $D,M]_{e}$ denotes the model that incorporate both the indexing map with multi-frame refiner and continuous indexing map estimator. As anticipated, the $D,M]_{e}$ configuration consistently yields the highest visual quality across all base models, demonstrating the effectiveness of both the multi-frame refiner and the continuous indexing map estimator.
Figure 12: Qualitative comparison under multi-frame setting. $[T]$: original arbitrary time VFI models using time indexing. $D]_{u}$: models trained using distance indexing, then inference using uniform maps. $D,R]_{u}$: models trained using both distance indexing and iterative reference-based estimation, then inference using uniform maps. $D,M]_{e}$: models trained using distance indexing with multi-frame fusion, then inference using estimated maps. Due to space limitations, the two adjacent input frames are omitted in the visualization, while the actual input consists of four frames.
VI-H Multi-frame quantitative comparison
In Table IX, we present a quantitative performance comparison across various model variants. The configurations $[T]$, $D]_{u}$, and $[D]$ retain the same definitions as previously described. $D]_{e}$ denotes the model augmented with a continuous indexing map estimator, $[D,M]$ includes the proposed multi-frame fusion with the refiner model, and $D,M]_{e}$ incorporates both the refiner and the map estimator. $[T,M]$ denotes multi-frame fusion using the refiner model trained with the time indexing map. On pixel-level metrics such as PSNR and SSIM, the enhanced variant $D,M]_{e}$ consistently outperforms the base model $[T]$ as well as $D]_{u}$, which lacks a map estimator and relies on uniform indexing maps. The improvement from $D]_{u}$ to $D,M]_{e}$ highlights the effectiveness of integrating the continuous indexing map estimator. Meanwhile, performance gains from $[D]$ to $[D,M]$, and from $D]_{e}$ to $D,M]_{e}$, demonstrate the additional benefit provided by the multi-frame refiner. Moreover, the comparison between time-indexed ($[T,M]$) and distance-indexed ($_{e}, [D,M][D,M]$) multi-frame video frame interpolation models again shows that our proposed distance indexing yields better interpolation results.
TABLE IX: Multi-frame comparison on Vimeo90K septuplet dataset. $[M]$ denotes the multi-frame fusion. $\cdot]_{e}$ denotes inference with estimated indexing map. Other notations have the same meaning as in the previous experiments.
RIFE [14] $[T]$ $[T,M]$ $D]_{u} [$$D]_{e} [$$M]_{e} [D][D$] $[D,M]$
PSNR $\uparrow$ 28.22
28.84
27.55 28.25 28.34 29.20
@@ VI-H Multi-frame quantitative comparison
IFRNet [22] $[T]$ $[T,M]$ $D]_{u} [$$D]_{e} [$$M]_{e} [D][D$] $[D,M]$
PSNR $\uparrow$ 28.26
28.75
27.40 27.63 28.28 29.25
@@
TABLE X: Multi-frame ablation study
@@
VI-I Multi-frame tuning strategy
In this section, we evaluate different training strategies for multi-frame fusion using RIFE [14] as the baseline for ablation studies. All experiments follow the same setting of previous experiments on Vimeo90k. We find that jointly tuning both refiner and map estimator leads to instability and failure to converge. Thus, we split the tuning process in two stages. In the first stage, we first finetune the multi-frame refiner $\mathcal{F^{\prime}}$ alone using original VFI loss. In the second stage, there are two possible choices. Since jointly tuning with the map estimator remains unstable, we either finetune the flow estimator only to adapt it to VFI (referred to as FE), or finetune the VFI modules $\mathcal{F}$ and $\mathcal{F^{\prime}}$ to adapt them to the pretrained flow estimator (referred to as FE), as shown in Table X. The loss term $\mathcal{L}_{D}$ is calculated by $\mathcal{L}_{D}=\|D_{t}-D_{t}^{\text{RAFT}}\|_{2}^{2}$, denoting the L2 loss between the indexing maps from CPFlow and RAFT. The setting $\mathcal{L}_{V_{f}}$ corresponds to training only the VFI refiner with its original loss, while $\mathcal{L}_{V_{b+f}}$ involves jointly training both the base VFI model and the refiner with their original loss. Among the tested strategies, jointly tuning both the VFI model and the refiner under $\mathcal{L}_{V_{b+f}}$ achieves the best performance. This improvement is attributed to the model’s ability to adapt both components to the updated indexing map derived from the newly predicted optical flow.
TABLE XI: Costs comparison including runtime and number of parameters. We evaluate the computational overhead on an NVIDIA A100 GPU using images with a resolution of 448×256.
$[D]$ $[D,M]$ $D,M]_{e}$
@@ VI-I Multi-frame tuning strategy
VI-J Computational costs of the proposed framework
@@
The runtime and the number of parameters are reported in Table XI. The estimation process introduces an additional latency of approximately 0.04 seconds per inference, which is relatively low. The number of parameters for the map estimator is 10.2M. Overall, the computational overheads introduced by indexing map estimation are acceptable in practical scenarios.
@@
We also evaluate the computational overhead introduced by the multi-frame refiner as reported in Table XI. The refiner adds approximately 0.03 seconds per inference and requires an additional 10.3 million parameters, which is practical for deployment.
VI-K Limitations
While the proposed distance indexing strategy effectively stabilizes motion interpolation under large temporal gaps, it does not explicitly address content ambiguity caused by occlusion or severe lighting variations, where visual information is partially missing or altered. The proposed distance indexing relies on distance maps estimated from optical flow. In scenarios involving severe occlusion, curved motion, or significant lighting variations, inaccuracies in optical flow estimation may lead to imperfect distance maps, which can negatively affect the learning process. While our experiments show that distance indexing remains robust to moderate estimation errors, its performance may be limited when such errors become dominant. Nevertheless, we believe that future advances in optical flow estimation will further enhance the effectiveness of distance-indexed video frame interpolation.
VII Conclusion
@@ Acknowledgments
We thank Dorian Chan, Zhirong Wu, and Stephen Lin for their insightful feedback and advice. Our thanks also go to Vu An Tran for developing the web application.
Appendix A Proof of Eq. (4)
Eq. (4) in the paper can be rigorously proven if an L2 loss is used,
@@ Appendix B The Rationale for Solving Ambiguity
Figure 13: Analysis of ambiguity for the case of multiple frames.
@@
We claim the “correct” distance map is hard to estimate accurately from merely two frames since there are a wide range of possible velocities. If considering more neighboring frames like Xu et al. [59] (more observation information), it is possible to estimate an accurate distance map for pixel-wise aligned interpolation.
@@
The case of multi-frame inputs
Multi-frame inputs introduce additional motion priors, which help constrain possible trajectories but cannot fully resolve velocity ambiguity. To clarify this, here is a theoretical illustration demonstrating that as shown in Fig. 13. And even when more frames are available, the velocity profile along the trajectory remains under-determined.
@@
TABLE XII: Comparison on Vimeo90K triplet dataset. $[T]$ denotes the method trained with traditional fixed time indexing paradigm. $[D]$ denotes the distance indexing paradigm. $\cdot]_{u}$ denotes inference with uniform map as time indexes.
@@ Appendix C Additional Experiments and Analysis
While the benefits of our proposed disambiguation strategies are best demonstrated on arbitrary-time VFI models, they actually improve the performance of fixed-time models as well. Using RIFE [14] as a representative example, we extend our comparison to the fixed-time training paradigm, depicted in Fig. 14. The label $[T]$ RIFE (Tri) refers to the model trained on the triplet dataset from Vimeo90K [61] employing time indexing. Conversely, $[D]$ RIFE (Tri) indicates training on the same triplet dataset but utilizing our distance indexing approach. Both $[D]$ RIFE and $[D,R]$ RIFE models are trained on the septuplet dataset, consistent with our earlier comparison. Despite being trained on varied datasets, it is evident that the arbitrary time model outperforms the fixed time model. However, the efficacy of distance indexing appears restrained within the fixed-time training paradigm. This limitation stems from the fact that deriving distance representation solely from the middle frame yields a sparse distribution, making it challenging for the network to grasp the nuances of distance. We delve deeper into the quantitative analysis of these findings in Table XII. Compared to training arbitrary time models on the septuplet dataset, the advantages of distance indexing become notably decreased when training fixed time models on the triplet dataset.
Two channel scheme
Comparison results with the two channel scheme, i.e., $D_{t}$ with $\bm{x}$ and $\bm{y}$ directions, are shown in Table XIII. The observations are as follows: (1) $D_{x,y}]_{u}$ outperforms $D]_{u}$, which makes sense since it accounts for both speed and direction. (2) $D,R]_{u}$ performs better than $D_{x,y}]_{u}$. Our speculation is that $[R]$ benefits not only from addressing the directional ambiguity but also from reducing the prediction difficulty via divide-and-conquer. (3) $D_{x,y},R]_{u}$ does not exceed $D,R]_{u}$, showing that the iterative formulation is sufficient to resolve the directional ambiguity. $D_{x,y}]_{u}$ has the potential to learn different trajectories but cannot realize that potential since the trajectory distribution within a short time is not sufficiently diverse. However, $D_{x,y}]_{u}$ only involves increasing the input channels and does not need to run iteratively. Thus, $D_{x,y}]_{u}$ is a better choice for fast/lightweight frame interpolation.
TABLE XIII: Comparison with $D_{x,y}]$ scheme on Vimeo90K septuplet dataset. $D_{x,y}]$ denotes training with two channel distance map with $\bm{x}$ and $\bm{y}$ directions. $\cdot]_{u}$ denotes inference with uniform maps.
$[T]$ $D]_{u} [$$D,R]_{u} [$$D_{x,y}]_{u} [$$D_{x,y},R]_{u}$
@@
Figure 15: Comparison on Vimeo90K Septuplet with GiMM-VFI [10].
TABLE XIV: Comparison on Vimeo90K Septuplet with GiMM-VFI [10] and BiM-VFI [41].
BiM-VFI GiMM-VFI $[T]$ GiMM-VFI $[D]$ GiMM-VFI $D]_{u}$
[…]
29.36
27.58
[…]
0.943
0.919
[…]
0.041
0.045
[…]
5.998
Evaluation on strong mainstream baselines
We further consider recent strong models based on mainstream VFI architectures, including GIMM-VFI [10] and BiM-VFI [41]. Since BiM-VFI can be viewed as a direct extension of our line of work, which introduces a new motion description (Bidirectional Motion Fields) built upon distance maps, we focus our evaluation on GIMM-VFI for a representative comparison. As shown in Table XIV, training GIMM-VFI with distance indexing consistently improves performance over its original time-indexed version. We also provide a qualitative comparison in Fig. 15. These results demonstrate that the proposed strategy remains effective even when applied to strong, modern architectures.
TABLE XV: Comparison on Adobe240 [47] for $\times$8 interpolation with RIFE [14].
$[T]$ $D]_{u} [$$D,R]_{u}$
PSNR $\uparrow$ 30.24
30.47
30.30
SSIM $\uparrow$
0.939
0.938 0.937
LPIPS $\downarrow$ 0.073 0.057
0.054
NIQE $\downarrow$ 5.206 4.974
4.907
TABLE XVI: Comparison on X4K1000FPS [44] for $\times$8 interpolation with RIFE [14].
$[T]$ $D]_{u} [$$D,R]_{u}$
PSNR $\uparrow$ 36.36
36.80
36.26
SSIM $\uparrow$
0.967
0.964 0.964
LPIPS $\downarrow$ 0.040 0.032
0.032
NIQE $\downarrow$ 7.130 6.936
6.924
More results on other benchmarks
Table XV and Table XVI show the results of RIFE [14] on Adobe240 [47] and X4K1000FPS [44] for $\times$8 interpolation respectively, using uniform maps. Distance indexing $D]_{u}$ and iterative reference-based estimation $R]_{u}$ strategies can consistently help improve the perceptual quality. In addition, it is noteworthy that $D]_{u}$ is better than $[T]$ in terms of the pixel-centric metrics like PSNR, showing that the constant speed assumption (uniform distance maps) holds well on these two easier benchmarks.
TABLE XVII: Performance comparison across motion-pattern categories for representative VFI models. The evaluation set is partitioned into four motion types based on optical-flow-derived velocity statistics: acceleration, deceleration, approximately constant-speed, and complex mixed acceleration/deceleration. Within each model and motion type, we compare uniform time indexing $[T]$ and ground-truth distance indexing $[D]$. The better result between $[T]$ and $[D]$ is highlighted in bold.
Model Motion Type Cat. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ NIQE $\downarrow$
RIFE [14] Acceleration $[T]$ 28.2145 0.9117 0.1015 6.7908
$[D]$ 29.2395 0.9290 0.0899 6.5830
Complex $[T]$ 28.4378 0.9126 0.1041 6.5606
$[D]$ 29.3960 0.9292 0.0912 6.3613
Constant $[T]$ 29.4517 0.9235 0.1017 6.1813
$[D]$ 30.2031 0.9347 0.0848 5.9581
Deceleration $[T]$ 27.8605 0.9088 0.1042 6.6731
$[D]$ 28.8007 0.9262 0.0971 6.5364
IFRNet [22] Acceleration $[T]$ 28.2330 0.9146 0.0869 6.5339
$[D]$ 29.2812 0.9314 0.0785 6.4588
Complex $[T]$ 28.4757 0.9161 0.0883 6.3191
$[D]$ 29.4851 0.9322 0.0782 6.2284
Constant $[T]$ 29.4212 0.9259 0.0850 6.0017
$[D]$ 30.2535 0.9370 0.0716 5.8317
Deceleration $[T]$ 27.8972 0.9123 0.0900 6.4507
$[D]$ 28.8555 0.9296 0.0841 6.3932
AMT-S [26] Acceleration $[T]$ 28.4748 0.9190 0.1008 6.9927
$[D]$ 29.6034 0.9368 0.0855 6.7747
Complex $[T]$ 28.7413 0.9211 0.1013 6.7628
$[D]$ 29.8756 0.9386 0.0832 6.5250
Constant $[T]$ 29.9110 0.9346 0.0934 6.4211
$[D]$ 30.7599 0.9457 0.0734 6.0756
Deceleration $[T]$ 28.1143 0.9160 0.1033 6.8830
$[D]$ 29.1998 0.9349 0.0917 6.7289
EMA-VFI [65] Acceleration $[T]$ 29.3060 0.9263 0.0871 6.8722
$[D]$ 30.2554 0.9410 0.0781 6.6792
Complex $[T]$ 29.7016 0.9302 0.0839 6.6009
$[D]$ 30.5960 0.9437 0.0746 6.4067
Constant $[T]$ 31.0516 0.9437 0.0708 6.1619
$[D]$ 31.6974 0.9511 0.0619 5.9517
Deceleration $[T]$ 28.9662 0.9245 0.0909 6.7890
$[D]$ 29.8220 0.9395 0.0838 6.6091
Figure 16: Qualitative comparison on acceleration motion patterns. Distance indexing ($[D]$) better preserves motion details and reduces blur artifacts compared to time indexing ($[T]$), consistent with the systematic results in Table XVII.
Figure 17: Qualitative comparison on deceleration motion patterns. Distance indexing ($[D]$) better preserves motion details and reduces blur artifacts compared to time indexing ($[T]$), consistent with the systematic results in Table XVII.
Systematic analysis on motion patterns
We provide a more direct and systematic analysis to quantify how distance indexing differentiates motion patterns here. Specifically, we categorize samples in the evaluation dataset of Vimeo90K Septuplet into four motion-pattern groups based on the temporal variation of optical flow magnitude (i.e., velocity statistics): acceleration, deceleration, approximately constant-speed, and complex mixed acceleration/deceleration. For each sample, we compute the optical flow magnitude for the intermediate frames and analyze its temporal trend. We then assign it to one of the above motion categories according to whether the velocity profile is increasing, decreasing, nearly constant, or exhibits mixed non-monotonic changes. We evaluate representative VFI models under uniform time indexing ($[T]$) and “ground-truth” distance indexing ($[D]$) within each motion category. The goal of this analysis is to verify whether distance indexing can consistently improve interpolation quality under diverse velocity profiles.
As shown in Table XVII, using accurate distance indexing consistently improves performance across all motion-pattern groups and across all evaluated models. Notably, the gains remain stable not only for approximately constant-speed motion, but also for accelerating and decelerating cases where velocity changes over time. This suggests that distance indexing provides a more expressive and physically meaningful motion parameterization, enabling the interpolated results to better align with diverse motion patterns.
In addition to the quantitative comparisons, we further provide representative qualitative examples for the acceleration and deceleration cases in Figs. 16 and 17, where the improvements in motion behavior and interpolation quality are visually consistent with the quantitative findings.
Appendix D Implementation Details
For VFI models that explicitly take the temporal index $t$ as input, we directly replace $t$ with the corresponding distance index $D_{t}$, without modifying the network architecture. For models that implicitly target a fixed interpolation time (e.g., predicting $I_{0.5}$), we concatenate the distance map $D_{t}$ with the RGB input frames as an additional channel. In practice, this only requires expanding the input channels of the first encoder layer, while keeping all network modules unchanged. For models with a shared encoder, the distance map is concatenated consistently with both input frames.
The same strategy applies to the iterative reference-based estimation setting ($[R]$), where the reference RGB image and distance map are concatenated to replace the original time indexing input. All changes are confined to the input representation and do not involve architectural redesign, but only a minimal expansion of the input channels.
Figure 18: User interface of user study.
Appendix E User Study UI